Okay I am so not getting this!!! I need help with solving by the susstitution method 4m + n = 33
m - 6n = 27
Any help at all will be greatly appreciated, I have a few other questions also, but can send them seperately. Thank you in advance, Blaine

Answer by rfer(16322)

m=6n+27
4m+n=33
4(6n+27)+n=33
24n+108+n=33
25n=-75
n=-3
-----------------
4m-3=33
4m=36
m=9
--------------
9-6*-3=27
9+18=27